BestBoiKirby:

Simone claims that the slope of the line through (-2,7) and (3,0) is the same as the slope of the line through (2,1) and (12,-13) prove or disprove simone's claim This is one of the questions on my homework, and I do not understand it at all. :(

BestBoiKirby:

@dude

BestBoiKirby:

@redhead23

BestBoiKirby:

@GirlBesty

Jeremana72:

idk

dude:

Do you know how to find the slope between two points?

redhead23:

yeah i am here to answer

BestBoiKirby:

only when its on a table. Idk how to make this a table tho.

BestBoiKirby:

so im pooped

redhead23:

so all u need to now is how to put it on the table?

BestBoiKirby:

but im also not rlly sure if i get it even with the table, im just hoping i do :(

redhead23:

ight i will help with that to well try to anyways

dude:

So, label a point as "point 1" and "point 2" For the first set you have (-2,7) and (3,0) (-2,7) is your "point 1" (3,0) is your "point 2" Remember that each point is written as (x,y) First subtract the y values (Second y point - first y point)

redhead23:

slope = (Y2-Y1)/(X2-X1) (0-7)/(3+2) = -7/5 (-13 -1) / (12-2) = -14/10 = -7/5 sure enough

BestBoiKirby:

so the y is zero and the x is one?

dude:

Careful, subtract only the y points Point 2 (y) minus "point 1" (y)

BestBoiKirby:

Sorry y is sevem

dude:

@redhead23 Yeah, do you understand what \(y_\#\) and \(x_\#\) are?

dude:

Yeah 0 - 7

BestBoiKirby:

ok, so what next?

dude:

You get -7 Now subtract the x values (Second x point minus first x point) Point 1: (-2,7) Point 2: (3,0)

BestBoiKirby:

so 1? or -5?

BestBoiKirby:

-5!

redhead23:

well isnt x 2 and y 1 i am not sure tho

redhead23:

wait i am wrong

BestBoiKirby:

no bc u r subtracting 3 from -2

BestBoiKirby:

so -5

dude:

Close Point 1: (-2,7) <- first point x is -2 Point 2: (3,0) <- second point x is 3 Second x minus first would be: \(3 - (- 2)\)

BestBoiKirby:

so 1?

dude:

Double negative is just positive \(3+2\)

BestBoiKirby:

OH RIGHT

BestBoiKirby:

5

dude:

Now divide the y subtracted and the x subtracted (y subtracted: -7) Earlier (x subtracted: 5) What we just did \(\dfrac{-7}{5}\) or just \(-\dfrac75\)

dude:

@redhead23 You were right but make sure you know what you are doing

dude:

We dont give out answers

redhead23:

oh sorry so i cant just say the answer i have to go through it and explain it to the person?

BestBoiKirby:

im confused what do i do after that

BestBoiKirby:

its a fraction, so do i do the same to the other side?

dude:

That's it

dude:

Now do the same for the other two points

redhead23:

ight my fault bro

BestBoiKirby:

and then see if the fractions are equal?

dude:

Yes

BestBoiKirby:

also i didnt see it redhead so ur good

BestBoiKirby:

Thank you dude? im gonna give u a medal, but can u give one to redhead as well?

BestBoiKirby:

thank you both!

redhead23:

well thank dude bc he went an explained it to u soo but yw tho

dude:

So basically 1. Subtract second y point from the first point 2. Subtract the second x point from the first 3. Divide what you got for the y over the x ~~~ If you're feeling confident the equation can help you \(\dfrac{y_2-y_1}{x_2-x_1}\) The number at the bottom just means which point it comes from Point 1: \((x_1,y_1)\) Point 2: \((x_2,y_2)\)